OpenStudy (firejay5):

I need the points from the vertices and the min and max. The ANSWERS ONLY Please! :D 22. Function: 5x - 2y y < or = to 7 - x 3x - 2y < or = to 6 2x + y > or = to -3

OpenStudy (mathlegend):

y is greater than or equal to 1 \[y \ge 1\] 3(1) + x 3(2)+ x 3(3) + x etc.... 2 is less than or equal to x x is less than or equal to 4 2\[2\le x \le 4\] So this number can be less than or equal to 4 4, 3, 2, 1, 0, -1, -2, -3, -4 etc... \[y \le -1/2x + 2\] -.50 + 2 = 1.5 so basically less than or equal to 1.5 1.4, 1.3, 1.2, 1.1, 1, 0, etc....

OpenStudy (firejay5):

@MathLegend Ya there

OpenStudy (firejay5):

I just need the coordinate points and if the equations have a minimum or a maximum

OpenStudy (anonymous):

find the intersection of all those lines, then check the function at those numbers

OpenStudy (firejay5):

do you know the answers just asking???

OpenStudy (firejay5):

can ya give me the answers, I know how to do it

OpenStudy (anonymous):

i don't now how to do it quickly, no the intersection of \(x=4\) and \(y=-\frac{1}{2}x+2\) is \((4,0)\) so you have to check that one

OpenStudy (firejay5):

what goes with what

OpenStudy (anonymous):

when \(x=2\) you get the point \((2,1)\) so you have to check that one as well

OpenStudy (anonymous):

first one is the \(x\) second is the \(y\)

OpenStudy (firejay5):

Is it 18, 20, or 22

OpenStudy (anonymous):

so you use \(3y + x\) at each point at \((4,0)\) you get \(3\times 0+4=4\) at \((2,1)\) you get \(4\times 1+2=6\)

OpenStudy (firejay5):

Are there are 2 points for #18

OpenStudy (anonymous):

wait i made a mistake somewhere

OpenStudy (anonymous):

oh it is \(3\times 1+2=5\) for the point \((2,1)\) that is the maximum

OpenStudy (anonymous):

these are a royal pain if you just want the answers, cheat (use technology)

OpenStudy (anonymous):

let me see if this works, hold on

OpenStudy (firejay5):

It's better to get it from you @satellite73

OpenStudy (anonymous):

nice huh note the syntax, finish quickly

OpenStudy (firejay5):

How do you make the > < or = to symbol on the website @satellite73

OpenStudy (anonymous):

\leq

OpenStudy (firejay5):

@satellite73 on number 20 and 22

OpenStudy (anonymous):